# Course Allocation with Credits via Stable Matching

### v1

@@ 5.1 No lower quotas
As we saw in the previous section, every instance of ca-ml has at least one stable matching, for any definition of stability. However, there may be more than one possible stable matching, and indeed they can have different sizes, in which case we want to find a stable matching that is as large as possible. We refer to this problem as ca-x-ml-max, where x is a definition of stability. ca-p-ml-max is trivially solvable in polynomial time, by Observation 3.
However, under the other three definitions of stability, ca-x-ml-max is NP-hard. In fact, under coalition stability, ca-c-ml-max is NP-hard even if courses and students are ordered in two master lists, every course has $1$ or $2$ credits and upper quota $1$, and every student can take up to $2$ credits.
In order to prove this result, we will transform from the Stable Marriage problem with Ties and Incomplete preferences (smti). com-smti, the problem of determining whether a complete weakly stable matching in an smti instance exists, is NP-complete even if there are master lists on both sides, the ties occur in one master list only, and are of length $2$ (irving_stable_2008, , Thm. 3.2).
The key to understanding the transformation from Theorem 5, which shows that ca-c-ml-max is NP-hard, is in Figure 5. In this instance, $c_{2}$ is a $2$-credit course and $c_{1}$ and $c_{3}$ are $1$-credit courses. Thus, under coalition stability, $s_{1}$ is indifferent between being matched with $c_{1}$ and $c_{3}$, and with $c_{2}$. Hence, under coalition stability, we can speak of a tie between $c_{2}$ on the one hand, and $c_{1}$ and $c_{3}$ on the other, even though preference lists are all strict. This idea gives the connection between com-smti and ca-c-ml-max.
@@
It is NP-hard to determine whether a ca-ml instance has a student-complete (or course-complete) coalition-stable matching, even if both students and courses are ordered according to master lists, every student takes up to $2$ credits, and every course is of size $1$ or $2$ with upper quota $1$. Therefore, ca-c-ml-max is NP-hard, even if there are no downward-feasible constraints.
@@
We transform from the following restriction of SMTI: in the SMTI instance $I$, both men and women are ordered in two master lists, ties only occur in the women’s master list (so on men’s preference lists only), and each tie is of length $2$. Under these conditions, COM-SMTI is NP-complete. We construct a ca-ml instance $J$ from $I$.
Every woman $w_{j}$ becomes a course $c_{j}$, with the same preference list, and with upper quota $1$. If $w_{j}$ is not involved in a tie in the women’s master list then $c_{j}$ is a $2$-credit course. If women $w_{k},w_{l}$ are tied in the women’s master list, then one of them (say $w_{k}$) becomes a $1$-credit course ($c_{k}$), while the other ($w_{l}$) becomes a $2$-credit course ($c_{l}$). For every $1$-credit course $c_{k}$ we create another $1$-credit course $c_{k}^{\prime}$. This course has the same preference list as $c_{k}$, substituting men for students, and it also has upper quota $1$. The master list of courses is the same as the master list of women, except that if women $w_{k},w_{l}$ are tied in the women’s master list, then the master list of courses (MLC) includes $c_{k}\succ_{MLC}c_{l}\succ_{MLC}c_{k}^{\prime}$, where $O(c_{k})=O(c_{k}^{\prime})=1$ and $O(c_{l})=2$.
Every man $m_{i}$ becomes a student $s_{i}$, who can take up to $2$ credits. The preferences of $s_{i}$ are inherited from $m_{i}$, such that every woman $w_{j}$ not involved in a tie in the women’s master list becomes $c_{j}$ in $s_{i}$’s preference list. If some $w_{j}$ is in a tie with $w_{k}$ in the women’s master list, but only $w_{j}$ is in the preference list of $m_{i}$, then we replace $w_{j}$ by $c_{j}$, if $c_{j}$ is a $2$-credit course, or by $c_{j}$ first and then $c_{j}^{\prime}$, if $c_{j}$ is a $1$-credit course. If both $w_{j}$ and $w_{k}$ are tied and in $m_{i}$’s preference list, then we substitute them both by $c_{j}\succ_{s_{i}}c_{k}\succ_{s_{i}}c_{j}^{\prime}$. By Figure 5, we know that $s_{i}$ is indifferent between being matched with $\{c_{j},c_{j}^{\prime}\}$ and with $\{c_{k}\}$, under coalition stability. From the men’s master list we derive a students’ master list. We now show that course-complete matchings are student-complete, and vice versa.
@@
Let $n$ be the number of men (and women) in the SMTI instance $I$. Then, in the ca-ml instance $J$, for each woman $w_{j}$ we have either one $2$-credit course with upper quota $1$, or two $1$-credit courses with upper quota $1$. Thus, a course-complete matching has a total of $2\cdot n$ credits. Similarly, for each man we have a student $s_{i}$ with $T(s_{i})=2$. Thus, a student-complete matching has $2\cdot n$ credits. Therefore, a course-complete matching has the same number of credits as a student-complete matching. ∎
@@
A consequence of this transformation is that, in a coalition-stable matching $M^{\prime}$, a student $s_{i}$ is matched with either a $2$-credit course $c_{j}$ or two $1$-credit courses $\{c_{k},c_{k}^{\prime}\}$, such that both courses come from the same woman $w_{k}$. Assume that this is not the case, and that there exists some $s_{i}$ that is matched with a $1$-credit course $c_{j}$ but not with $c_{j}^{\prime}$ (the other case, where $s_{i}$ is matched to $c_{j}^{\prime}$ instead, can be proved analogously). Furthermore, assume that $s_{i}$ is the highest placed student in this situation and that $c_{j}$ is ranked higher than the other course matched to $s_{i}$, if there is one. Then, because $s_{i}$ is the highest ranked student matched to one but not both of the courses that come from the same woman, $c_{j}^{\prime}$ prefers $s_{i}$ over its only student. Similarly, if $s_{i}$ is matched to some $1$-credit course $c_{l}$, then she prefers $c_{j}^{\prime}$ over $c_{l}$. Therefore, $(s_{i},\{c_{j}^{\prime}\})$ is a blocking coalition, a contradiction to the stability of $M^{\prime}$. Hence, if a student $s_{i}$ is matched with a $1$-credit course $c_{j}$, she is also matched with $c_{j}^{\prime}$.
@@
Now we show that there exists a complete weakly stable matching in $I$ if and only if there exists a student-complete (course-complete) coalition-stable matching in $J$. First, assume that there is a complete weakly stable matching $M$ in $I$. Then we build a complete matching $M^{\prime}$ in $J$ as follows. For each $(m_{i},w_{j})\in M$ such that $c_{j}$ is a $2$-credit course, add $(s_{i},c_{j})$ to $M^{\prime}$. For each $(m_{i},w_{j})\in M$ such that $c_{j}$ is a $1$-credit course, add $(s_{i},c_{j})$ and $(s_{i},c_{j}^{\prime})$ to $M^{\prime}$. Clearly the matching is both student-complete and course-complete, as every course is matched, and every student takes $2$ credits. We now show that $M^{\prime}$ is coalition-stable. Assume that it is not, and that there exists a blocking coalition $(s_{p},\{c_{q}\})$, with $O(c_{q})=2$ (or $(s_{p},\{c_{q},c_{q}^{\prime}\})$ where $O(c_{q})=1$). Thus, $s_{p}$ strictly prefers $c_{q}$ (and $c_{q}^{\prime}$, if it exists) over every course in $C_{M^{\prime}}(s_{p})$. This implies that $c_{q}$ (and $c_{q}^{\prime}$, if it exists) is placed higher in the courses’ master list than every course in $C_{M^{\prime}}(s_{p})$. Similarly, $c_{q}$ (and $c_{q}^{\prime}$, if it exists) strictly prefers $s_{p}$ over its student $s_{r}$, meaning that $s_{p}$ is placed higher in the students’ master list than $s_{r}$. However, this means than in $I$, $w_{q}$ prefers $m_{p}$ over $m_{r}$, and that $m_{p}$ prefers $w_{q}$ over his partner in $M$, meaning that $(m_{p},w_{q})$ blocks $M$, a contradiction to the stability of $M$. Thus, $M^{\prime}$ is a coalition-stable matching.
@@
Now, assume that $J$ has a student-complete (course-complete) coalition-stable matching $M^{\prime}$. Thus, every student $s_{i}$ is matched with either a $2$-credit course $c_{j}$, or two $1$-credit courses, $c_{k}$ and $c_{k}^{\prime}$. We build a complete matching $M$ in $I$ as follows. If $(s_{i},c_{j})\in M^{\prime}$, add $(m_{i},w_{j})$ to $M$. Clearly $M$ is complete, because every man and every woman is matched. We now show that $M$ is weakly stable. Assume that it is not, so that there exists a blocking pair $(m_{p},w_{q})$. Thus, $m_{p}$ prefers $w_{q}$ over his partner in $M$, and $w_{q}$ prefers $m_{p}$ over her partner, which means that $w_{q}$ and $m_{p}$ are ranked higher in their respective master lists than their partners. This implies that $c_{q}$ (or $c_{q}$ and $c_{q}^{\prime}$) and $s_{p}$ are ranked higher in their respective master lists than the courses in $C_{M^{\prime}}(s_{p})$ and the only student in $S_{M^{\prime}}(c_{q})$, respectively. Hence, $s_{p}$ strictly prefers $c_{q}$ (and $c_{q}^{\prime}$, if it exists) over every course in $C_{M^{\prime}}(s_{p})$. As $c_{q}$ (and $c_{q}^{\prime}$) also strictly prefers $s_{p}$ over its only assigned student, $(s_{p},\{c_{q}\})$ (or $(s_{p},\{c_{q},c_{q}^{\prime}\})$) is a blocking coalition, which contradicts the stability of $M^{\prime}$. Thus, $M$ is a weakly stable matching.
@@
Likewise, determining whether a ca-ml instance has a course-complete first-coalition-stable matching is NP-hard, as well as determining whether it has a course-complete pair-size-stable matching. Hence, both ca-fc-ml-max and ca-ps-ml-max are NP-hard. However, in the following proof, we assume that any two courses have a different number of credits.
@@
It is NP-hard to determine whether there exist a course-complete first-coalition-stable matching, even if both students and courses are ordered according to master lists, and every course has upper quota $1$. Hence, both ca-fc-ml-max and ca-ps-ml-max are NP-hard, even if there are no downward-feasible constraints.
@@
We reduce from COM-SMTI. Recall that COM-SMTI is NP-complete even if there are master lists on both sides, the ties occur in one master list only (in this case the women’s list), and are of length $2$. The ca-ml instance is $J$.
@@
Every woman $w_{j}$ becomes a course $c_{j}$, with the same preference list, and with upper quota $1$. The first course in the master list, $c_{1}$, has number of credits $O(c_{1})=1+n\cdot\varepsilon$, where $n$ is the number of courses and $\varepsilon$ is an arbitrarily small positive number. The second course, $c_{2}$, has $O(c_{2})=1+(n-1)\cdot\varepsilon$, and so on, so that the $i$-th course, $c_{i}$, has number of credits $O(c_{i})=1+(n-i+1)\cdot\varepsilon$. If two courses, $c_{j}$ and $c_{j+1}$, are tied, one (say $c_{j}$) has number of credits $O(c_{j})=1+(n-j+1)\cdot\varepsilon$ and the other, $c_{j+1}$, has number of credits $O(c_{j+1})=1+(n-j)\cdot\varepsilon$. A property of this construction is that courses that are higher in the master list have a larger number of credits than those placed below. The master list of courses (MLC) mostly comes from the women’s master list. However, if two women $w_{j}$ and $w_{j+1}$ are tied in the master list, then the course with the smallest number of credits (say $c_{j+1}$) is ranked before the course with the larger number of credits (say $c_{j}$). That is, if $O(c_{j+1})<O(c_{j})$, then $c_{j+1}\succ_{MLC}c_{j}$. This way, a student matched with $c_{j}$ cannot form a first-blocking coalition with $c_{j+1}$, and vice versa.
@@
Every man $m_{i}$ becomes a student $s_{i}$, with $T(s_{i})=1+n\cdot\varepsilon$. The preferences of $s_{i}$ are inherited from $m_{i}$, so that if $w_{j}$ is in $m_{i}$’s preference list, $c_{j}$ is in $s_{i}$’s preference list. The order of the courses in given by the master list of courses. It is easy to see that no student has enough credits for more than one course.
@@
First, assume that there is a complete weakly stable matching $M$ in $I$. Then we build a course-complete matching $M^{\prime}$ in $J$ as follows. If $(m_{i},w_{j})\in M$, add $(s_{i},c_{j})$ to $M^{\prime}$. The matching is course-complete, as every course is matched with one student. We now show that $M^{\prime}$ is first-coalition-stable. Assume that it is not, and that there exists a first-blocking coalition $(s_{p},\{c_{q}\})$. Thus, $s_{p}$ strictly prefers $c_{q}$ over the course in $C_{M^{\prime}}(s_{p})$, and $O(c_{q})>O(C_{M^{\prime}}(s_{p}))$. This implies that $c_{q}$ is placed higher in the courses’ master list than the course in $C_{M^{\prime}}(s_{p})$. Similarly, $c_{q}$ strictly prefers $s_{p}$ over its only matched student $s_{r}$, meaning that $s_{p}$ is placed higher in the students’ master list than $s_{r}$. However, this means than in $I$ $w_{q}$ prefers $m_{p}$ over $m_{r}$, and that $m_{p}$ prefers $w_{q}$ over his partner, meaning that $(m_{p},w_{q})$ blocks $M$, a contradiction to the stability of $M$. Thus, $M^{\prime}$ is a first-coalition-stable matching.
@@
Now, assume that $J$ has a course-complete first-coalition-stable matching $M^{\prime}$. Thus, every student $s_{i}$ is matched with one course, because no student has enough credits to be matched with two courses. We build a complete matching $M$ in $I$ as follows. For each $(s_{i},c_{j})\in M^{\prime}$, add $(m_{i},w_{j})$ to $M$. The matching is complete, as every man is matched with one woman. We now show that $M$ is stable, too. Assume that it is not, so that there exists a blocking pair $(m_{p},w_{q})$. Thus, $m_{p}$ prefers $w_{q}$ over his partner, and $w_{q}$ prefers $m_{p}$ over her partner, which means that $w_{q}$ and $m_{p}$ are ranked higher in their respective master lists than their respective partners. This implies that $c_{q}$ and $s_{p}$ are ranked higher in their respective master lists than the course in $C_{M^{\prime}}(s_{p})$ and the student in $S_{M^{\prime}}(c_{q})$, respectively. Hence, $s_{p}$ strictly prefers $c_{q}$ over the course in $C_{M^{\prime}}(s_{p})$, and $O(c_{q})>O(C_{M^{\prime}}(s_{p}))$. As $c_{q}$ also strictly prefers $s_{p}$ over its only assigned student, $(s_{p},\{c_{q}\})$ is a first-blocking coalition, which contradicts the stability of $M^{\prime}$. Thus, $M$ is a weakly stable matching.
@@
In fact, the hardness of finding a maximum size first-coalition-stable matching remains even if every course has $1$ or $2$ credits. In order to prove this result, we will reduce from Min-MM. In this problem, we have a graph $G$ and a positive integer $K$, and the goal is to determine whether there exists a maximal matching of size at most $K$. This problem is NP-complete, even if $G$ is a subdivision graph horton_1993_minimum , which implies that it is bipartite and all vertices in one of the sides have degree $2$.
@@
It is NP-hard to determine whether an instance of ca-ml has a course-complete first-coalition-stable matching, even if every student takes up to $2$ credits, every course has $1$ or $2$ credits, and each student has up to $5$ courses in her preference list.

### v2

@@ 5.1 No lower quotas
As we saw in Section 4.2, every instance of ca-ml has at least one stable matching, for any definition of stability. However, there may be more than one possible stable matching, and indeed they can have different sizes, in which case a natural goal is to find a stable matching that is as large as possible. We refer to this problem as ca-x-ml-max, where x is a definition of stability. ca-p-ml-max is trivially solvable in polynomial time, by Observation 3.
However, under the other three definitions of stability, we will show in this section that ca-x-ml-max is NP-hard. We will begin by showing that, under coalition stability, ca-c-ml-max is NP-hard even if courses and students are ordered in two master lists, every course has one or two credits and upper quota $1$, and every student can take up to two credits.
In order to prove this result, we will transform from the stable marriage with ties and incomplete preferences problem (smti) Manlove et al. (2002). com-smti, the problem of determining whether a complete weakly stable matching in an smti instance exists, is NP-complete even if there are master lists on both sides, the ties occur in one master list only, and are of length $2$ (Irving et al., 2008, Thm. 3.2).
Theorem 5, which shows that ca-c-ml-max is NP-hard, involves a key idea that is illustrated in Figure 5. In this instance, $c_{2}$ is a $2$-credit course and $c_{1}$ and $c_{3}$ are $1$-credit courses. Thus, under coalition stability, $s_{1}$ is indifferent between being matched with $c_{1}$ and $c_{3}$, and with $c_{2}$. Hence, under coalition stability, we can speak of a tie between $c_{2}$ on the one hand, and $c_{1}$ and $c_{3}$ on the other, even though preference lists are all strict. This idea gives the connection between com-smti and ca-c-ml-max.
@@
It is NP-hard to determine whether a ca-ml instance has a student-complete (or course-complete) coalition-stable matching, even if both students and courses are ordered according to master lists, every student takes up to two credits, and every course has one or two credits and upper quota $1$. Therefore, ca-c-ml-max is NP-hard, even if there are no downward-feasible constraints.
@@
We transform from the restriction of com-smti in which both men women are ordered in two master lists, ties only occur in the women’s master list (so on men’s preference lists only), and each tie is of length $2$. Under these conditions, com-smti is NP-complete Irving et al. (2008). Let $I$ be an instance of this restricted version of com-smti. We construct a ca-ml instance $J$ from $I$.
Every woman $w_{j}$ becomes a course $c_{j}$, with the same preference list, and with upper quota $1$. If $w_{j}$ is not involved in a tie in the women’s master list then $c_{j}$ is a $2$-credit course. If women $w_{k},w_{l}$ are tied in the women’s master list, then one of them (say $w_{k}$) becomes a $1$-credit course ($c_{k}$), while the other ($w_{l}$) becomes a $2$-credit course ($c_{l}$). For every $1$-credit course $c_{k}$ we create another $1$-credit course $c_{k}^{\prime}$. This course initially has the same preference list as $c_{k}$, and it also has upper quota $1$. The master list of courses is the same as the master list of women, except that if women $w_{k},w_{l}$ are tied in the women’s master list, then the master list of courses ($\succ_{MLC}$) includes $c_{k}\succ_{MLC}c_{l}\succ_{MLC}c_{k}^{\prime}$, where $O(c_{k})=O(c_{k}^{\prime})=1$ and $O(c_{l})=2$.
Every man $m_{i}$ becomes a student $s_{i}$, who can take up to two credits. On the preference list of every course in $J$, replace any man $m_{i}$ by student $s_{i}$. The preferences of $s_{i}$ are inherited from $m_{i}$, such that every woman $w_{j}$ not involved in a tie in the women’s master list becomes $c_{j}$ in $s_{i}$’s preference list. If some $w_{j}$ is in a tie with $w_{k}$ in the women’s master list, but only $w_{j}$ is in the preference list of $m_{i}$, then we replace $w_{j}$ by $c_{j}$, if $c_{j}$ is a $2$-credit course, or by $c_{j}$ first and then $c_{j}^{\prime}$, if $c_{j}$ is a $1$-credit course. If both $w_{j}$ and $w_{k}$ are tied and in $m_{i}$’s preference list, then we substitute them both by $c_{j}\succ_{s_{i}}c_{k}\succ_{s_{i}}c_{j}^{\prime}$. By Figure 5, we know that $s_{i}$ is indifferent between being matched with $\{c_{j},c_{j}^{\prime}\}$ and with $\{c_{k}\}$, under coalition stability. From the men’s master list we derive a students’ master list ($\succ_{MLS}$). We now show that course-complete matchings are student-complete, and vice versa.
@@
Let $n$ be the number of men (and women) in the com-smti instance $I$. Then, in the ca-ml instance $J$, for each woman $w_{j}$ we have either one $2$-credit course with upper quota $1$, or two $1$-credit courses with upper quota $1$. Thus, a course-complete matching has a total of $2n$ credits. Similarly, for each man we have a student $s_{i}$ with $T(s_{i})=2$. Thus, a student-complete matching has $2n$ credits. Therefore, a course-complete matching has the same number of credits as a student-complete matching. ∎
We next show that in a coalition-stable matching $M^{\prime}$ in $J$, a student $s_{i}$ is matched with either a $2$-credit course or two $1$-credit courses that come from the same woman in $I$.
Lemma 2.
Let $M^{\prime}$ be a coalition-stable matching in $J$. Then any student $s_{i}$ is matched in $M^{\prime}$ with either a $2$-credit course $c_{j}$ or two $1$-credit courses $\{c_{k},c_{k}^{\prime}\}$, such that both courses come from the same woman $w_{k}$ in $I$.
Proof.
@@
Assume that this is not the case, and that there exists some $s_{i}$ that is matched in $M^{\prime}$ with a $1$-credit course $c_{j}$ but not with $c_{j}^{\prime}$ (the other case, where $s_{i}$ is matched to $c_{j}^{\prime}$ instead, can be proved analogously). Furthermore, assume that $s_{i}$ is the highest placed student in $\succ_{MLS}$ in this situation, and without loss of generality assume that $c_{j}$ is ranked higher in $\succ_{MLC}$ than the other course matched to $s_{i}$, if there is one. Then, because $s_{i}$ is the highest ranked student in $\succ_{MLS}$ matched to one but not both of the courses that come from the same woman, $c_{j}^{\prime}$ is unmatched or prefers $s_{i}$ over its only student. Similarly, if $s_{i}$ is matched to some $1$-credit course $c_{l}$, then she prefers $c_{j}^{\prime}$ over $c_{l}$. Therefore, $(s_{i},\{c_{j}^{\prime}\})$ is a blocking coalition, a contradiction to the stability of $M^{\prime}$. Hence, if a student $s_{i}$ is matched with a $1$-credit course $c_{j}$ in $M^{\prime}$, she is also matched with $c_{j}^{\prime}$ in $M^{\prime}$. ∎
@@
Now we show that there exists a complete weakly stable matching in $I$ if and only if there exists a student-complete (or course-complete by Lemma 1) coalition-stable matching in $J$. First, assume that there is a complete weakly stable matching $M$ in $I$. Then we build a complete matching $M^{\prime}$ in $J$ as follows. For each $(m_{i},w_{j})\in M$ such that $c_{j}$ is a $2$-credit course, add $(s_{i},c_{j})$ to $M^{\prime}$. For each $(m_{i},w_{j})\in M$ such that $c_{j}$ is a $1$-credit course, add $(s_{i},c_{j})$ and $(s_{i},c_{j}^{\prime})$ to $M^{\prime}$. Clearly the matching is both student-complete and course-complete, as every course is matched, and every student takes two credits. We now show that $M^{\prime}$ is coalition-stable. Assume that it is not, and that there exists a blocking coalition $(s_{p},\{c_{q}\})$, with $O(c_{q})=2$ (or $(s_{p},\{c_{q},c_{q}^{\prime}\})$ where $O(c_{q})=1$). Thus, $s_{p}$ prefers $c_{q}$ (and $c_{q}^{\prime}$, if it exists) over every course in $C_{M^{\prime}}(s_{p})$. This implies that $c_{q}$ (and $c_{q}^{\prime}$, if it exists) is placed higher in $\succ_{MLC}$ than every course in $C_{M^{\prime}}(s_{p})$. Similarly, $c_{q}$ (and $c_{q}^{\prime}$, if it exists) prefers $s_{p}$ over its student $s_{r}$, meaning that $s_{p}$ is placed higher in $\succ_{MLS}$ than $s_{r}$. However, this means than in $I$, $w_{q}$ prefers $m_{p}$ over $m_{r}$, and that $m_{p}$ prefers $w_{q}$ over his partner in $M$, meaning that $(m_{p},w_{q})$ blocks $M$, a contradiction to the weak stability of $M$ in $I$. Thus, $M^{\prime}$ is a coalition-stable matching in $J$.
@@
Now, assume that $J$ has a student-complete (course-complete) coalition-stable matching $M^{\prime}$. Then, by Lemma 2, every student $s_{i}$ is matched with either a $2$-credit course $c_{j}$, or two $1$-credit courses, $c_{k}$ and $c_{k}^{\prime}$. We build a complete matching $M$ in $I$ as follows. If $(s_{i},c_{j})\in M^{\prime}$, add $(m_{i},w_{j})$ to $M$. Clearly $M$ is complete, because every man and every woman is matched. We now show that $M$ is weakly stable. Assume that it is not, so that there exists a blocking pair $(m_{p},w_{q})$. Thus, $m_{p}$ prefers $w_{q}$ over his partner in $M$, and $w_{q}$ prefers $m_{p}$ over her partner in $M$, which means that $w_{q}$ and $m_{p}$ are ranked higher in their respective master lists than their partners. This implies that $c_{q}$ (or $c_{q}$ and $c_{q}^{\prime}$) and $s_{p}$ are ranked higher in their respective master lists than the courses in $C_{M^{\prime}}(s_{p})$ and the only student in $S_{M^{\prime}}(c_{q})$, respectively. Hence, $s_{p}$ prefers $c_{q}$ (and $c_{q}^{\prime}$, if it exists) over every course in $C_{M^{\prime}}(s_{p})$. As $c_{q}$ (and $c_{q}^{\prime}$) also prefers $s_{p}$ over its only assigned student, $(s_{p},\{c_{q}\})$ (or $(s_{p},\{c_{q},c_{q}^{\prime}\})$) is a blocking coalition of $M^{\prime}$, which contradicts the coalition stability of $M^{\prime}$ in $J$. Thus, $M$ is a weakly stable matching in $I$.
@@
Likewise, determining whether a ca-ml instance has a course-complete first-coalition-stable matching is NP-hard, as well as determining whether it has a course-complete pair-size-stable matching, as we will now show. Hence, both ca-fc-ml-max and ca-ps-ml-max are NP-hard.
@@
It is NP-hard to determine whether there exists a course-complete first-coalition-stable matching, even if both students and courses are ordered according to master lists, and every course has upper quota $1$. The same result also holds for pair-size stability. Hence, both ca-fc-ml-max and ca-ps-ml-max are NP-hard, even if there are no downward-feasible constraints.
@@
We again transform from the NP-complete restriction of com-smti in which both men and women are ordered in two master lists, ties only occur in the women’s master list, and each tie is of length $2$ Irving et al. (2008). Let $I$ be an instance of this restricted version com-smti. We construct a ca-ml instance $J$ from $I$.
@@
Every woman $w_{j}$ becomes a course $c_{j}$ with upper quota $1$. The master list of courses $\succ_{MLC}$ in $J$ is obtained from the master list of women in $I$ by replacing every woman $w_{j}$ by course $c_{j}$. We will ensure that any two courses have a different number of credits, as follows. Assume that the courses in $\succ_{MLC}$ are ordered $c_{1},c_{2},\dots,c_{n}$ such that $c_{1}\succeq_{MLC}c_{2}\succeq_{MLC}\dots c_{n-1}\succeq_{MLC}c_{n}$, where $n$ is the number of courses, and $c_{r}\succeq_{MLC}c_{s}$ denotes that $c_{r}$ comes before $c_{s}$ on $\succ_{MLC}$ or the two courses are tied in $\succ_{MLC}$. For each $j$ ($1\leq j\leq n$), let $c_{j}$ have number of credits $O(c_{j})=1+(n-j+1)\varepsilon$, where $\varepsilon$ is an arbitrarily small positive number (e.g., $\varepsilon=\frac{1}{n}$). Next, if two courses $c_{j}$ and $c_{j+1}$ are tied in $\succ_{MLC}$, swap $O(c_{j})$ and $O(c_{j+1})$. This ensures that if $c_{r}\succ_{MLC}c_{s}$ then $O(c_{r})>O(c_{s})$, whilst if $c_{r}$ and $c_{s}$ are tied in $\succ_{MLC}$ then $O(c_{r})<O(c_{s})$. Hence a student matched with $c_{s}$ cannot form a first-blocking coalition with $c_{r}$, and vice versa.
@@
Every man $m_{i}$ becomes a student $s_{i}$, with $T(s_{i})=1+n\cdot\varepsilon$. The master list of students $\succ_{MLS}$ in $J$ is obtained from the master list of men in $I$ by replacing every man $m_{i}$ by student $s_{i}$. It is easy to see that no student has enough available credits for more than one course.
@@
First, assume that there is a complete weakly stable matching $M$ in $I$. Then we build a course-complete matching $M^{\prime}$ in $J$ as follows. If $(m_{i},w_{j})\in M$, add $(s_{i},c_{j})$ to $M^{\prime}$. The matching is course-complete, as every course is matched with one student. We now show that $M^{\prime}$ is first-coalition-stable in $J$. Assume that it is not, and that there exists a first-blocking coalition $(s_{p},\{c_{q}\})$. Thus, $s_{p}$ prefers $c_{q}$ over the course in $C_{M^{\prime}}(s_{p})$, and $O(c_{q})\geq O(C_{M^{\prime}}(s_{p}))$. This implies that $c_{q}$ is placed higher in $\succ_{MLC}$ than the course in $C_{M^{\prime}}(s_{p})$. Similarly, $c_{q}$ prefers $s_{p}$ over its only matched student $s_{r}$, meaning that $s_{p}$ is placed higher in $\succ_{MLS}$ than $s_{r}$. However, this means than in $I$, $w_{q}$ prefers $m_{p}$ over $m_{r}$, and that $m_{p}$ prefers $w_{q}$ over his partner, meaning that $(m_{p},w_{q})$ blocks $M$, a contradiction to the weak stability of $M$ in $I$. Thus, $M^{\prime}$ is a first-coalition-stable matching in $J$.
@@
Now, assume that $J$ has a course-complete first-coalition-stable matching $M^{\prime}$. Thus, every student $s_{i}$ is matched with one course, because no student has enough available credits to be matched with two courses. We build a complete matching $M$ in $I$ as follows. For each $(s_{i},c_{j})\in M^{\prime}$, add $(m_{i},w_{j})$ to $M$. The matching is complete, as every man is matched with one woman. We now show that $M$ is weakly stable in $I$, too. Assume that it is not, so that there exists a blocking pair $(m_{p},w_{q})$. Thus, $m_{p}$ prefers $w_{q}$ over his partner, and $w_{q}$ prefers $m_{p}$ over her partner, which means that $w_{q}$ and $m_{p}$ are ranked higher in their respective master lists than their respective partners. This implies that $c_{q}$ and $s_{p}$ are ranked higher in their respective master lists than the course in $C_{M^{\prime}}(s_{p})$ and the student in $S_{M^{\prime}}(c_{q})$, respectively. Hence, $s_{p}$ prefers $c_{q}$ over the course in $C_{M^{\prime}}(s_{p})$, and $O(c_{q})>O(C_{M^{\prime}}(s_{p}))$. As $c_{q}$ also prefers $s_{p}$ over its only assigned student, $(s_{p},\{c_{q}\})$ is a first-blocking coalition, which contradicts the first-coalition stability of $M^{\prime}$ in $J$. Thus, $M$ is a weakly stable matching in $I$.
@@
In fact, the problem of finding a maximum size first-coalition-stable matching in ca-ml remains NP-hard even if every course has one or two credits. In order to prove this result, we will reduce from min-mm. In this problem, we have a graph $G$ and a positive integer $K$, and the goal is to determine whether there exists a maximal matching of size at most $K$. This problem is NP-complete, even if $G$ is a subdivision graph Horton and Kilakos (1993), which implies that it is bipartite and all vertices in one of the sides have degree $2$.
@@
It is NP-hard to determine whether an instance of ca-ml has a course-complete first-coalition-stable matching, even if every student takes up to two credits, every course has one or two credits, and each student has up to $5$ courses in her preference list.